# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@
Abstract: This is a collection of articles, written as sections, on arithmetic properties of differential equations, holomorphic foliations, Gauss-Manin connections and Hodge loci. Each section is independent from the others and it has its own abstract and introduction and the reader might get an insight to the text by reading the introduction of each section. The main connection between them is through comments in footnotes. Our major aim is to develop a theory of leaf schemes over finitely generated subrings of complex numbers, such that the leaves are also equipped with a scheme structure. We also aim to formulate a local-global conjecture for leaf schemes.
Introduction
The analogies and relations between complex and arithmetic algebraic geometry have occupied the attention of mathematicians for more than a century. The most concrete example of this is Deligne’s theory of mixed Hodge structures in [Del71] mainly inspired by results in $p$-adic cohomologies. In the same time, linear differential equations under the modern name of connections on vector bundles, have been investigated as a branch of algebraic geometry, and such analogies and relations have been enlarged to this context. The most famous example is the Grothendieck-Katz $p$-curvature conjecture for algebraicity of solutions of linear differential equations and local systems. It was suggested by Grothendieck and popularized by Katz in [Kat70, Kat72]. Whereas Deligne’s theory is based upon analogies between cohomology theories, and consequently there is no concrete bridge between complex and $p$-adic Hodge theories, Grothendieck-Katz conjecture is a concrete statement about the modulo primes properties of linear differential equations and its consequence on their solutions over complex numbers. One of the main goals of the present text is to put the Grothendieck-Katz conjecture into a larger framework of leaf schemes and Hodge loci and investigate a local-global principle for such objects. Our study aims to provide a new approach to the Cattani-Deligne-Kaplan theorem in [CDK95] on the algebraicity of Hodge loci by a local analysis of Taylor series of periods, whereas the original proof uses the global analysis of the monodromy group. We aim to introduce modulo primes tools for proving some conjectures in [Mov21a, Chapter 19] and [Mov21b]. The first one predicts the existance of certain Hodge loci for cubic hypersurfaces of dimension $6$ and $8$ and the second one provides a conjectural counter example to a conjecture of J. Harris for special components of Noether-Lefschetz loci in the case of octic surfaces.
Beyond linear differential equations, we have the theory of holomorphic foliations in complex manifolds which has its origin in the second part of Hilbert’s 16th problem on the uniform boundedness of number of limit cycles. In the last decades, it has also been investigated as a branch of algebraic geometry. Despite this, considering foliations on non-reduced schemes or with non-reduced leaves might seem abstract non-sense. In the present text we develop a theory in which a holomorphic foliation is identified with its module of differential $1$-forms (which may not be saturated) and so it is not just the underlying geometric object. Moreover, its leaves might be non-reduced with different codimenions for which we use the name leaf scheme. The main examples of leaf schemes are Hodge loci. The origin of the theory of leaf schemes goes back to [Mov17b] and [Mov22, Chapter 5,6] and we follow and further develop these texts. For experts in holomorphic foliations the theory of holomorphic foliations in which leaves are replaced with leaf schemes might seem an abstract non-sense. This is mainly due to the lack of motivation and simple examples. We are mainly motivated by applications to Hodge loci, even though the corresponding foliations cannot be written explicitly.
Summary of the text:
Each section is independent from the others and it has its own abstract and introduction. The best way to get an insight to the text is by reading the introduction of each section. In order to achieve this we have allowed ourselves to repeat definitions and notations if this makes the reading smoother. The connection between sections is through comments in footnotes. The content of Section 3 is the highlight of the present text. In this section we investigate a local-global principle for leaf schemes. Natural examples of leaf schemes come from the study of Hodge loci. A conjectural description of some Hodge loci is presented in Section 5 and the content of this section is the goal behind our main goals. The local-global principle studied here in the case of linear differential equations or connections on vector bundles is known as the Grothendieck-Katz p-curvature conjecture, and in Section 1 we rise some doubts about this conjecture. For this reason in Section 2 we analyse this conjecture for Lamé equations. In Section 1 we also analyse a local-global conjecture for vector fields. In Section 4 we discuss arithmetic properties of the Ramanujan vector field.
Acknowledgement:
Since the first messy draft of the present text was written, it was in the author’s webpage for sharing ideas. Throughout the text, there appears the name of many mathematicians who have made comments regarding parts of the text. Email, online and personal communications with them have improved the text a lot. The reader after a PDF search of their names can find the place of their contributions. This includes: João Pedro dos Santos, Jorge V. Pereira, Yves André, Hélène Esnault, Wadim Zudilin, Stefan Reiter, Frits Beukers, Daniel Litt, Joshua Lam, Nick Katz, David Urbanik, Felipe Voloch, Frederico Bianchini, Pierre Deligne, Remke Kloosterman, Roberto Villaflor, Florian Fürnsinn. My sincere thanks go to all of them.
1 Grothendieck-Katz conjecture
Man sollte weniger danach streben, die Grenzen der mathematischen Wissenschaften zu erweitern, als vielmehr danach, den bereits vorhandenen Stoff aus umfassenderen Gesichtspunkten zu betrachten, E. Study, (see the preface of [Edw90]).
@@ 1.1 Introduction
(1) $\frac{\partial y}{\partial z}={\sf A}(z)y,$
@@
[…]
@@
If ${\mathfrak{R}}$ is a ring of integers of a number field, in the above theorem we can take $N\in\mathbb{N}$, see 1.2. This is natural as in the left hand side of (2) only integers are inverted and not an element of ${\mathfrak{R}}-\mathbb{Z}$. The only reason that we assume that ${\mathfrak{R}}$ is the ring of integers of a number field is 1.5. The conclusion in the second part of 1.1 is equivalent to $\Delta^{m}{\sf A}_{m}\equiv_{p^{k}}0$ which means $\Delta^{m}{\sf A}_{m}=0$ in the ring ${\mathfrak{R}}[z,\frac{1}{N}]/p^{k}{\mathfrak{R}}[z,\frac{1}{N}]$. We may expect that the converse of 1.1 holds.
@@
For a linear differential equation (1) if there is $N\in{\mathfrak{R}}$ such that for all $m\in\mathbb{N}$ the entries of $\frac{(N\Delta)^{m}{\sf A}_{m}}{m!}$ are in ${\mathfrak{R}}[z]$ then its solutions are algebraic functions over $\mathbb{C}(z)$.
@@
For a prime number $p$, the matrix ${\sf A}_{p}$ considered as a matrix with entries in ${\mathfrak{R}}[z,\frac{1}{\Delta}]/p{\mathfrak{R}}[z,\frac{1}{\Delta}]$ is usually called the $p$-curvature of ${\sf A}$ (there is no $N$ in the denominator). The statement in the second part of 1.1 for $k=1$ and $m=p$ and with $N\in\mathbb{N}$ (see 1.2) says that the $p$-curvature vanishes for all except a finite number of primes. Its converse, which is a stronger version of 1.1, is known as Grothendieck-Katz (p-curvature) conjecture. It was suggested by A. Grothendieck and popularized by N. Katz in [Kat70, Kat72]. Note that there is a finite number of primes which are invertible in ${\mathfrak{R}}[\frac{1}{N}]$ and the second statement in the main theorem becomes empty for these primes. Therefore, the Grothendieck-Katz conjecture is: If for all but a finite number of primes the $p$-curvature of $y^{\prime}={\sf A}y$ vanishes then all its solutions are algebraic.
@@
Therefore, any proof of the conclusion of 1.1 without assuming the Grothendieck-Katz conjecture, must use a mixed prime argument. João Pedro dos Santos took my attention to the articles [MvdP03] in which the authors right at the end of the article makes the following conjecture: Let $F$ be a number field and $M$ be a differential module over $F(z)$. If for all but a finite number of primes $p$, the iterative differential module with respect to $p$ exists and has a finite differential Galois group $G$, then the differential Galois group of $M$ is isomorphic to $G$. The property (2) implies the existence of such an iterative differential module and so 1.1 implies this conjecture. A similar conjecture has been stated in [EK18] using the language of $D$-modules: Let $X/\mathbb{C}$ be a smooth projective variety over $\mathbb{C}$ and $M=(E,\nabla)$ be a vector bundle with an integrable connection on $X$. If there is a dense open subscheme $U\subset{\rm Spec}({\mathfrak{R}})$ such that for all closed points $s\in U$, $M_{s}$ underlies a $D_{X_{s}}$-module then $M$ has finite monodromy. Another similar conjecture has been advertised by M. Kontsevich in [Kon]. Hopefully the reader will be convinced that the hypothesis of conjectures in [MvdP03, EK18, Kon] are hard to check whereas (2) can be even implemented in a computer and it is characteristic zero statement. Moreover, we would like to emphasize that more divisibility properties are missing in the hypothesis of Grothendieck-Katz conjecture. Y. André in a personal communication took my attention to the works [BS82, DGS94, And04]. The following is written based on his comments. The $p$-adic radius of convergence of a full set of solutions of a linear differential equation is $\geq p^{-1/(p-1)}$, see [BS82], and if this is strict, for instance if the $p$-curvature is nilpotent, then $\frac{{\sf A}_{m}(z)}{m!}$ becomes divisible by very high powers of $p$ for large $m$. The vanishing of $p$-curvatures for a fixed $p$ implies that the generic radius $R_{p}$ of convergence satisfies $R_{p}>p^{-1/(p-1)}$ and the $\tau$-invariant is 00 in this case, see [And04, Section 5]. H. Esnault informed me about [EG20, Proposition 8.1, 8.2] in which the authors verify the Grothendieck-Katz conjecture for two partial cases. W. Zudilin has called phenomenon like (2) the cancellation of factorials, see [Zud01]. He took my attention to [DR22] in which the interest in algebraic solutions of rank one systems is attributed to Abel.
@@ 1.2 Algebraic functions
The following proposition for ${\mathfrak{R}}=\mathbb{Z}$ is due to G. Eisenstein, see [Lan04] and the references therein. In the following $y(z)$ is a holomorphic function $y:U\to\mathbb{C}$ defined in some connected open subset $U$ of $\mathbb{C}$.
@@
$\Delta y_{n}=\hbox{ a polynomial of degree $\leq n$ in $y_{i},\ 1\leq i<n$, with coefficients in ${\mathfrak{R}}$},$
for instance $\Delta y_{1}=-\frac{\partial P}{\partial z}(z_{0},y_{0})\in{\mathfrak{R}}$. This implies that $y_{n}$ lies in the finitely generated $\mathbb{Z}$-algebra ${\mathfrak{R}}[\frac{1}{\Delta}]$. More precisely, by induction we can show that $y_{n}$ has a pole order at most $12n-$ 1 at $\Delta$. For $n=1$ this is easy. If this is true for all $m<n$ then $\Delta y_{n}$ is a sum of monomials $y_{i_{1}}y_{i_{2}}\cdots y_{i_{k}}$ with $i_{1}+i_{2}+\cdots+i_{k}\leq n$ and $1\leq i_{1},i_{2},\ldots,i_{k}<n$ and coefficients in ${\mathfrak{R}}$. If $k=1$ then $22i_{1}-1\leq 2n-$ 2 and we are done. If $k\geq 2$ then $22i_{1}-1+2i_{2}-1\cdots+2i_{k}-1\leq 2n-k\leq 2n-$ 2 and we are done again. For the second part of the proposition we put $N:=\Delta^{2}$. ∎
@@
Let $\check{\mathfrak{R}}$ be a finitely generated $\mathbb{Z}$-subalgebra of $\mathbb{C}$ containing the coefficients ${\sf A}$, $z_{0}\in\mathbb{C}$, $\frac{1}{\Delta(z_{0})}$ and the entries of $y_{0}=[y_{0,1},y_{0,2},\cdots,y_{0,n}]^{{\mathsf{t}}{\mathsf{r}}}\in\mathbb{C}^{\sf n}$. A solution of $y^{\prime}={\sf A}y$ with the initial condition $y(z_{0})=y_{0}$ is a formal power series with coefficients in $\check{\mathfrak{R}}_{\mathbb{Q}}:=\check{\mathfrak{R}}\otimes_{\mathbb{Z}}\mathbb{Q}$:
@@
where we have written the Taylor series ${\sf A}=\sum_{i=0}^{\infty}{\sf A}_{i}(z-z_{0})^{i}$. Since the entries of ${\sf A}$ are rational functions with coefficients in $\check{\mathfrak{R}}$, the entries of ${\sf A}_{i}$ are in $\check{\mathfrak{R}}$. This follows from the fact that for a polynomial $P\in\check{\mathfrak{R}}[z]$ we have:
@@
Let ${\mathfrak{R}}$ be as Section 1.1, that is, ${\mathfrak{R}}$ is a finitely generated $\mathbb{Z}$-subalgebra of $\mathbb{C}$ containing the coefficients of $\Delta$ and the coefficients of the polynomials in $\Delta(z){\sf A}$. By 1.4 we have a ${\sf n}\times{\sf n}$ matrix $Y(z)$ whose entries are formal power series in $z-z_{0}$ and with coefficients in the ring ${\mathfrak{R}}_{\mathbb{Q}}[z_{0},\frac{1}{\Delta(z_{0})}]$, $Y(z_{0})$ is the identity matrix and $dY={\sf A}Y$. This is called the fundamental system at $z=z_{0}$. From another side we know that the entries of $Y$ are algebraic, and so by 1.1 we can assume that the corresponding polynomials are defined over ${\mathfrak{R}}_{\mathbb{Q}}[z_{0},\frac{1}{\Delta(z_{0})}]$. A multiplication by a power of $\Delta(z_{0})$ and possibly an integer, we can assume that such polynomials are defined over ${\mathfrak{R}}[z_{0}]$. Now by 1.3 we know that the coefficients of the Taylor series of $Y$ are in ˇ$\check{\mathfrak{R}}:={\mathfrak{R}}[z_{0},\frac{1}{N}]$ for some $N\in{\mathfrak{R}}[z_{0}]$. This $N$ is the product of all $N$’s attached to the polynomial equation of each entry of $Y$. Since for any $f=(z-z_{0})^{n}$ we have
$f^{(m)}=\binom{n}{m}m!(z-z_{0})^{n-m},$
we conclude that $Y^{(m)}=0$ in (ˇ$(\check{\mathfrak{R}}/m!\check{\mathfrak{R}})[[z-z_{0}]]$, and hence
(3) ${\sf A}_{m}Y=0\hbox{ in }(\check{\mathfrak{R}}/m!\check{\mathfrak{R}})[[z-z_{0}]].$
From now on take $z_{0}\in{\mathfrak{R}}$ and hence ˇ$\check{\mathfrak{R}}={\mathfrak{R}}[\frac{1}{N}]$. The number $N$ in the announcement of theorem is exactly the number $N$ that we have obtained from Eisenstein theorem. We have ˇ$\check{\mathfrak{R}}/m!\check{\mathfrak{R}}=0$ if and only if
(4) $m!|N^{d},\hbox{ in ${\mathfrak{R}}$ for some $d\in\mathbb{N}$}.$
In this case (2) is automatic. Therefore, we assume that ˇ$\check{\mathfrak{R}}/m!\check{\mathfrak{R}}\not=0$, and in particular 10101$\not$=01 ≠ 0 in this ring. In this ring $Y=I_{{\sf n}\times{\sf n}}+\sum_{i=1}^{\infty}y_{i}(z-z_{0})^{i}$ is invertible, that is the entries of $Y^{-1}$ are formal power series in $(z-z_{0})$ and coefficients in ˇ$\check{\mathfrak{R}}/m!\check{\mathfrak{R}}$. This together with (3) imply that the Taylor series of ${\sf A}_{m}(z)$, and hence $\Delta(z)^{m}{\sf A}_{m}(z)$, at $z=z_{0}$, is zero in ˇ$\check{\mathfrak{R}}/m!\check{\mathfrak{R}}[[z-z_{0}]]$. But $R(z):=\Delta(z)^{m}{\sf A}_{m}(z)$ is a matrix of polynomials, and hence, its entries are in $m!\check{\mathfrak{R}}[z]$. This finishes the proof of the fact that the entries of $\frac{\Delta^{m}\cdot{\sf A}_{m}}{m!}$ are ${\mathfrak{R}}[z,\frac{1}{N}]$, for all $m\in\mathbb{N}$ which is slightly weaker than (2). The proof of this stronger statement is similar. ∎

### v2

# Leaf schemes and Hodge loci [^1] December 14, 2025

@@ 1.1 Introduction
(1.1) $\frac{\partial y}{\partial z}={\sf A}(z)y,$
@@
(1.2)  $\hbox{ entries of }\frac{(N\Delta)^{m}\cdot{\sf A}_{m}}{m!}\ \in\ {\mathfrak{R}}[z],\ \ \forall m\in\mathbb{N}.$
@@
If $\mathfrak{R}$} is a ring of integers of a number field, in the above theorem we can take $N\in\mathbb{N}$, see 1.2. This is natural as in the left hand side of (1.2) only integers are inverted and not an element of $\mathfrak{R}$$\mathbb{Z}$. The only reason that we assume that $\mathfrak{R}$} is the ring of integers of a number field is 1.5. The conclusion in the second part of 1.1 is equivalent to $\Delta^{m}{\sf A}_{m}\equiv_{p^{k}}0$ which means $\Delta^{m}{\sf A}_{m}=0$ in the ring ${\mathfrak{R}}[z,\frac{1}{N}]/p^{k}{\mathfrak{R}}[z,\frac{1}{N}]$. We may expect that the converse of 1.1 holds.
@@
For a linear differential equation (1.1) if there is $N\in{\mathfrak{R}}$ such that for all $m\in\mathbb{N}$ the entries of $\frac{(N\Delta)^{m}{\sf A}_{m}}{m!}$ are in ${\mathfrak{R}}[z]$ then its solutions are algebraic functions over $\mathbb{C}(z)$.
@@
For a prime number $p$, the matrix ${\sf A}_{p}$ considered as a matrix with entries in ${\mathfrak{R}}[z,\frac{1}{\Delta}]/p{\mathfrak{R}}[z,\frac{1}{\Delta}]$ is usually called the $p$-curvature of $\sf A$} (there is no $N$ in the denominator). The statement in the second part of 1.1 for $k=1$ and $m=p$ and with $N\in\mathbb{N}$ (see 1.2) says that the $p$-curvature vanishes for all except a finite number of primes. Its converse, which is a stronger version of 1.1, is known as Grothendieck-Katz (p-curvature) conjecture. It was suggested by A. Grothendieck and popularized by N. Katz in [Katz1970, Katz1972]. Note that there is a finite number of primes which are invertible in ${\mathfrak{R}}[\frac{1}{N}]$ and the second statement in the main theorem becomes empty for these primes. Therefore, the Grothendieck-Katz conjecture is: If for all but a finite number of primes the $p$-curvature of $y^{\prime}={\sf A}y$ vanishes then all its solutions are algebraic.
@@
Therefore, any proof of the conclusion of 1.1 without assuming the Grothendieck-Katz conjecture, must use a mixed prime argument. João Pedro dos Santos took my attention to the articles [MatzatPut2003] in which the authors right at the end of the article makes the following conjecture: Let $F$ be a number field and $M$ be a differential module over $F(z)$. If for all but a finite number of primes $p$, the iterative differential module with respect to $p$ exists and has a finite differential Galois group $G$, then the differential Galois group of $M$ is isomorphic to $G$. The property (1.2) implies the existence of such an iterative differential module and so 1.1 implies this conjecture. A similar conjecture has been stated in [KisinEsnault2018] using the language of $D$-modules: Let $X/\mathbb{C}$ be a smooth projective variety over $\mathbb{C}$ and $M=(E,\nabla)$ be a vector bundle with an integrable connection on $X$. If there is a dense open subscheme $U\subset{\rm Spec}({\mathfrak{R}})$ such that for all closed points $s\in U$, $M_{s}$ underlies a $D_{X_{s}}$-module then $M$ has finite monodromy. Another similar conjecture has been advertised by M. Kontsevich in [Kontsevich2023]. Hopefully the reader will be convinced that the hypothesis of conjectures in [MatzatPut2003, KisinEsnault2018, Kontsevich2023] are hard to check whereas (1.2) can be even implemented in a computer and it is characteristic zero statement. Moreover, we would like to emphasize that more divisibility properties are missing in the hypothesis of Grothendieck-Katz conjecture. Y. André in a personal communication took my attention to the works [BombieriSperber1982, dw94, Andre2004]. The following is written based on his comments. The $p$-adic radius of convergence of a full set of solutions of a linear differential equation is $\geq p^{-1/(p-1)}$, see [BombieriSperber1982], and if this is strict, for instance if the $p$-curvature is nilpotent, then $\frac{{\sf A}_{m}(z)}{m!}$ becomes divisible by very high powers of $p$ for large $m$. The vanishing of $p$-curvatures for a fixed $p$ implies that the generic radius $R_{p}$ of convergence satisfies $R_{p}>p^{-1/(p-1)}$ and the $\tau$-invariant is 0 in this case, see [Andre2004, Section 5]. H. Esnault informed me about [Esnault2020, Proposition 8.1, 8.2] in which the authors verify the Grothendieck-Katz conjecture for two partial cases. W. Zudilin has called phenomenon like (1.2) the cancellation of factorials, see [Zudilin2001]. He took my attention to [Delaygue-Rivoal] in which the interest in algebraic solutions of rank one systems is attributed to Abel.
@@ 1.2 Algebraic functions
The following proposition for $\mathfrak{R}$$\mathbb{Z}$ is due to G. Eisenstein, see [Landau1904] and the references therein. In the following $y(z)$ is a holomorphic function $y:U\to\mathbb{C}$ defined in some connected open subset $U$ of $\mathbb{C}$.
@@
[…]
for instance $\Delta y_{1}=-\frac{\partial P}{\partial z}(z_{0},y_{0})\in{\mathfrak{R}}$. This implies that $y_{n}$ lies in the finitely generated $\mathbb{Z}$-algebra ${\mathfrak{R}}[\frac{1}{\Delta}]$. More precisely, by induction we can show that $y_{n}$ has a pole order at most 2n12n-1 at $\Delta$. For $n=1$ this is easy. If this is true for all $m<n$ then $\Delta y_{n}$ is a sum of monomials $y_{i_{1}}y_{i_{2}}\cdots y_{i_{k}}$ with $i_{1}+i_{2}+\cdots+i_{k}\leq n$ and $1\leq i_{1},i_{2},\ldots,i_{k}<n$ and coefficients in $\mathfrak{R}$}. If $k=1$ then $2i_{1}-1\leq 2n-2$ and we are done. If $k\geq 2$ then $2i_{1}-1+2i_{2}-1\cdots+2i_{k}-1\leq 2n-k\leq 2n-2$ and we are done again. For the second part of the proposition we put $N:=\Delta^{2}$. ∎
@@
[…]
@@
[…]
@@
Let $\mathfrak{R}$} be as Section 1.1, that is, $\mathfrak{R}$} is a finitely generated $\mathbb{Z}$-subalgebra of $\mathbb{C}$ containing the coefficients of $\Delta$ and the coefficients of the polynomials in $\Delta(z){\sf A}$. By 1.4 we have a ${\sf n}\times{\sf n}$ matrix $Y(z)$ whose entries are formal power series in $z-z_{0}$ and with coefficients in the ring ${\mathfrak{R}}_{\mathbb{Q}}[z_{0},\frac{1}{\Delta(z_{0})}]$, $Y(z_{0})$ is the identity matrix and $dY={\sf A}Y$. This is called the fundamental system at $z=z_{0}$. From another side we know that the entries of $Y$ are algebraic, and so by 1.1 we can assume that the corresponding polynomials are defined over ${\mathfrak{R}}_{\mathbb{Q}}[z_{0},\frac{1}{\Delta(z_{0})}]$. A multiplication by a power of $\Delta(z_{0})$ and possibly an integer, we can assume that such polynomials are defined over ${\mathfrak{R}}[z_{0}]$. Now by 1.3 we know that the coefficients of the Taylor series of $Y$ are in $\check{\mathfrak{R}}:={\mathfrak{R}}[z_{0},\frac{1}{N}]$ for some $N\in{\mathfrak{R}}[z_{0}]$. This $N$ is the product of all $N$’s attached to the polynomial equation of each entry of $Y$. Since for any $f=(z-z_{0})^{n}$ we have
$f^{(m)}=\binom{n}{m}m!(z-z_{0})^{n-m},\$
we conclude that $Y^{(m)}=0$ in $(\check{\mathfrak{R}}/m!\check{\mathfrak{R}})[[z-z_{0}]]$, and hence
(1.3) ${\sf A}_{m}Y=0\hbox{ in }(\check{\mathfrak{R}}/m!\check{\mathfrak{R}})[[z-z_{0}]].$
From now on take $z_{0}\in{\mathfrak{R}}$ and hence $\check{\mathfrak{R}}={\mathfrak{R}}[\frac{1}{N}]$. The number $N$ in the announcement of theorem is exactly the number $N$ that we have obtained from Eisenstein theorem. We have $\check{\mathfrak{R}}/m!\check{\mathfrak{R}}=0$ if and only if
(1.4) $m!|N^{d},\hbox{ in ${\mathfrak{R}}$ for some $d\in\mathbb{N}$}.$
In this case (1.2) is automatic. Therefore, we assume that $\check{\mathfrak{R}}/m!\check{\mathfrak{R}}\not=0$, and in particular $1\not=0$ in this ring. In this ring $Y=I_{{\sf n}\times{\sf n}}+\sum_{i=1}^{\infty}y_{i}(z-z_{0})^{i}$ is invertible, that is the entries of $Y^{-1}$ are formal power series in $(z-z_{0})$ and coefficients in $\check{\mathfrak{R}}/m!\check{\mathfrak{R}}$. This together with (1.3) imply that the Taylor series of ${\sf A}_{m}(z)$, and hence $\Delta(z)^{m}{\sf A}_{m}(z)$, at $z=z_{0}$, is zero in $\check{\mathfrak{R}}/m!\check{\mathfrak{R}}[[z-z_{0}]]$. But $R(z):=\Delta(z)^{m}{\sf A}_{m}(z)$ is a matrix of polynomials, and hence, its entries are in $m!\check{\mathfrak{R}}[z]$. This finishes the proof of the fact that the entries of $\frac{\Delta^{m}\cdot{\sf A}_{m}}{m!}$ are ${\mathfrak{R}}[z,\frac{1}{N}]$, for all $m\in\mathbb{N}$ which is slightly weaker than (1.2). The proof of this stronger statement is similar. ∎